The price of blue socks is twice the price of black socks. If John buys 20 pairs of black socks and 30 pairs of blue socks he pays $120. How much does one pair of blue socks cost ?
(A) $1.50 (B) $3.00 (C) $4.50 (D) $1.00 (E) none of these

Answer by mananth(16949)

The price of blue socks is twice the price of black socks. If John buys 20 pairs of black socks and 30 pairs of blue socks he pays $120. How much does one pair of blue socks cost ?
(A) $1.50 (B) $3.00 (C) $4.50 (D) $1.00 (E) none of these
let cost of black socks be x
So blue socks will be 2x
20x+30*2x=120
20x+60x=120
80x=120
x=120/80
$1.5 black socks
$3 will be the cost of blue socks